OpenStudy (anonymous):

-6/7y<10 solve for y how do I do this?

OpenStudy (anonymous):

the -6/7 is a fraction

OpenStudy (anonymous):

assuming it is \[-\frac{6}{7}y<10\] you do it by multiplying both sides by \(-\frac{7}{6}\) and also changing the inequality, since \(-\frac{7}{6}\) is negative

OpenStudy (anonymous):

i.e. . \[-\frac{6}{7}y<10\iff y>-\frac{7}{6}\times 10\]

OpenStudy (anonymous):

how do you multiply -7/6 x 10?